OpenStudy (buttermequeasy):

If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4). 17 −17 5 −5

OpenStudy (buttermequeasy):

Please help. :c

OpenStudy (mertsj):

(f-g)(X) = 2x-2 -(3x-1)= 2x-3x-2+1=-x-1 (f-g)(4) - -4-1=-5

jhonyy9 (jhonyy9):

f(x)-g(x)=2x-2-3x+1=-x-1 (f-g)(4)=-4-1=-5

jhonyy9 (jhonyy9):

d is right